Express y in terms of x , it being given that 3x+y-12=0.Check whether the point (3,3) is on the line represented by the equation 3x+y-12=0.
Answer by rothauserc(4718)

3x + y -12 = 0
subtract 3x from both sides of =
y - 12 = -3x
add 12 to both sides of =
y = -3x + 12
the problem asks if the point (3,3) is on the line, substitute for x and y
3 = -3*3 +12
3 = -9 + 12
3 = 3
therefore the point (3,3) lies on the line y = -3x + 12